1. In assembling a nucleosome, normally the …(1) histone dimers first combine to form a tetramer, which then further combines with two … (2) histone dimers to form the octamer.

A. 1: H1–H3; 2: H2A–H2B

B. 1: H3–H4; 2: H2A–H2B

C. 1: H2A–H2B; 2: H1–H3

D. 1: H2A–H2B; 2: H3–H4

E.   1: H1–H2; 2: H3–H4

2. There has been a mutation on several enzymes important in DNA replication. For each enzyme, discuss what the protein normally does and how the mutation would impact DNA replication.   

a. The mutation affects editing site of the DNA polymerase III so that any incorrect nucleotide cannot slide into the editing site.

b. A mutation occurs which renders DNA ligase does not function.

c. A mutation occurs so that DNA primase does not function.

d. A mutation occurs so that the sliding clamp can no longer bind to the DNA strand.

3. If 30 percent of bases on a strand of DNA are thymine

a. what percentage of bases are adenine?

b. what percentage of bases are cytosine?

c. what percentage of bases are purines?

d. what percentage of bases are pyrimidines?
Answer #1

1.The correct answer would be:B. 1: H3–H4; 2: H2A–H2B

B. 1: H3–H4; 2: H2A–H2B

Reason:During the process of assembly of a nucleosome first H3–H4 and H2A-H2B dimers are formed and then the H3–H4 dimers combine to form tetramers. the H3–H4 tetramer then goes on to combines with two H2A-H2B dimers and this forms the compact octamer core.

2.

The mutation affects editing site of the DNA polymerase III so that any incorrect nucleotide cannot slide into the editing site :DNA polymerase III posseses the ability of proof reading in which any mismatched nucleotide /incorrect base pair is recognised by DNA polymerase III ,excised and then correct nucleotide is inserted,When mutation occurs in editing site of DNA polymerase III ,this incorrect nucleotide will not be correct and the DNA replicarion will continue with the error.

b. A mutation occurs which renders DNA ligase does not function :DNA ligase functions in filling the nicks in the leading strand and in re-annealing the semi-conservative strands and joining the Okazaki Fragments of the lagging strand. When mutation will occur the reannealing of the nicks in leading strand and the okazaki fragments of lagging strand will not be possible.

c. A mutation occurs so that DNA primase does not function:Primase enzyme functions to add primers to the template strands so that DNA replication can start.When the enzyme gets mutated, the primers will not be available for DNA replication to start and DNA replication will not start.

d. A mutation occurs so that the sliding clamp can no longer bind to the DNA strand: Sliding clamp with clamp proteins is formed by clamp proteins that attach to DNA so that DNA polymerase can make contact with the DNA template and carry out the processivity .When mutation occurs, DNA polymerase contact with template will be hindered ,affecting the elongation of replication process.

3.

As per chargaff's rule,

A+G= T+C

Adenine pairs with thymine and guanine pairs with cytosine

if 30% bases are thymine,

then A pairs with T .So, A will be 30%

Now, A+ T = 30+ 30= 60%

so, G+ C = 100 - 60 = 40%

again G pairs with C,

So each will be 20%

a. percentage of bases that are adenine =30%

b. what percentage of bases are cytosine =20%

c. what percentage of bases are purines (adenine + guanine) = 30+ 20 = 50%

d. what percentage of bases are pyrimidines(cytosine+thymine) = 20+30 =50%